From 4 distinct herbs and 3 distinct veggies, how many salads can be formed consisting of 3 different herbs and 2 different veggies?
Answer by checkley77(12844)

I appologize for using the wrong formula.
I used the permutation nPr but should have used the combination nCr formula.
4C3*3C2=4*3=12.